OpenStudy (anonymous):

if (3-4i)(x+iy) = 1+ i.0 find x , y belongs to R , i = √-1

OpenStudy (anonymous):

solve (3-4i)(x+iy)...you vl get wo terms...wih and without iota...equate the term without iota with 1 and with iota with 0...

OpenStudy (anonymous):

thank yu .. can you help me even better .. because when my teacher taught me i wer absent .. so i dont realy kno how to do this .. what is ( i ) ?

OpenStudy (anonymous):

i is iota...with the value of root -1...in dis question...iota trerm has to be equated to the rhs iota term and real term is equated with rhs real term...ask any odr question u vnt to ask..

OpenStudy (anonymous):

okay .. thanks .. :) so now i have to solve (3-4i)(x+iy) right ?

OpenStudy (anonymous):

yes...

OpenStudy (unklerhaukus):

\[(3-4i)(x+iy) = 1+0i\]\[3x+3yi-4ix-4yi^2=1+0i\]\[3x+(3y-4x)i+4y=1+0i\]\[(3x+4y)+(3y-4x)i=1+0i\] \[3x+4y=1;\qquad3y-4x=0\]

OpenStudy (anonymous):

thank you @UnkleRhaukus and @harsimran :)

OpenStudy (anonymous):

anytym :)